Hi Everybody! I've been out of school for 23 years and even when I was there I was never any good at this stuff. Any help would be Appreciated.
Here's the problem:
2[-7y-3(5-4y)+10]=10y-12

Answer by stanbon(75887)

2[-7y-3(5-4y)+10]=10y-12
Divide both sides by 2 to get:
[-7y-3(5-4y)+10] = 5y - 6
[-7y -15 + 12y + 10] = 5y -6
5y -5 = 5y -6
If you subtract 5y from both sides you get this contradiction:
-5 = -6
================
The equation is "inconsistent" and has no solution.
The thing that was inconsistent was all those "y's".
You thought they all meant the same number; they don't.
===================
Welcome back to the land of the learning.
------------------------
Cheers,
Stan H.